I am having trouble figuring how to set up and solve this equation:
Solve for x: (1/4) ^2x+8 = (16) ^ 2x+5
I have tried to guess numbers an put them in practice but I am not clear what operations I need to use to follow this equation
I appreciate your help!!

Answer by stanbon(75887)

Solve for x: (1/4) ^2x+8 = (16) ^ 2x+5
--------------
(2^-2)^(2x+8) = (2^4)^(2x+5)
2^(-4x-16) = 2^(8x+20)
Since the bases are the same the exponents are equal:
-4x-16 = 8x+20
12x = -36
x = -3
============
Cheers,
Stan H.